# Survival Analysis with Adversarial Regularization

### v3

@@ IV-B Adversarial Perturbations
Fast Gradient Sign Method (FGSM) [45] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [42] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation of $\boldsymbol{x}$ to remain within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
@@
$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),$ (16)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)},$ (17)
where $k$ is the iteration step and $\alpha$ is a step size parameter.
IV-C Random Noise
@@ V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing models robust to dataset uncertainty and variability, as motivated in Section I. We propose a method for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and stability on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with adversarial regularization as a Min-Max optimization problem (Eq. 12) [41, 40], as detailed in the following sections.
@@ V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [11]. Instead, the proposed SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of Eq. 10:
@@ V-B Adversarial Robustness Regularization
V-C SAWAR Training
The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [47], where we use early stopping [48] with respect to the validation set to prevent overfitting We introduce the covariate perturbations after 100 epochs. A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after 100 epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
@@ VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete. As evaluation metrics, we consider the Concordance Index (CI) [49], the Integrated Brier Score (IBS) [50] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1. To demonstrate that SAWAR produces SA models that perform well and are robust to dataset uncertainty, this section quantifies the improvement in predictive accuracy, calibration, and population curve estimation.
VI-A Configuration
The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively. The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions. We use lifelines [51] and sksurv [39] for computing SA metrics.
VI-B Datasets
@@
We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates.
@@ VI-C Adversary Setup
We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B) and a novel worst-case perturbation. FGSM is applied by using the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective Eq. 20, such that:
DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
1.00 2.9 3.35 3.35 2.5 2.9 4.6 4.4 2.4 1.9 1.7 4.0 4.7 2.7 2.1 1.5
0.90 3.1 3.5 3.45 2.45 2.5 4.5 4.5 2.5 1.8 1.7 4.0 4.7 2.9 1.9 1.5
0.80 3.4 3.4 3.2 2.35 2.65 4.6 4.4 2.6 1.8 1.6 4.1 4.6 3.0 1.9 1.4
0.70 3.45 3.35 3.1 2.4 2.7 4.7 4.3 2.6 1.8 1.6 4.2 4.5 3.0 2.0 1.3
0.60 3.45 3.45 3.0 2.5 2.6 4.7 4.3 2.7 1.8 1.5 4.2 4.5 3.1 2.1 1.1
0.50 3.85 3.65 3.1 2.1 2.3 4.4 4.6 2.9 1.8 1.3 4.1 4.5 3.2 2.1 1.1
0.40 3.9 3.8 3.05 2.1 2.15 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.30 4.4 4.2 3.0 1.9 1.5 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.20 4.35 4.0 3.3 2.25 1.1 4.2 4.6 3.2 2.0 1.0 3.9 4.7 3.2 2.1 1.1
0.10 4.25 4.0 3.3 2.35 1.1 4.0 4.6 3.2 2.2 1.0 3.8 4.8 3.1 2.3 1.0
0.05 3.85 4.0 3.45 2.6 1.1 3.8 4.7 2.9 2.2 1.4 3.9 4.7 3.1 2.3 1.0
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
1.00 2.8 3.4 3.5 3.6 1.7 2.6 4.3 3.8 3.3 1.0 3.1 4.5 3.4 3.0 1.0
0.90 3.1 3.3 3.5 3.4 1.7 3.0 4.3 3.6 3.1 1.0 3.1 4.5 3.5 2.9 1.0
0.80 3.0 3.3 3.5 3.7 1.5 3.0 4.5 3.5 3.0 1.0 3.0 4.6 3.5 2.9 1.0
0.70 3.05 3.3 3.6 3.65 1.4 3.0 4.5 3.5 3.0 1.0 3.1 4.5 3.5 2.9 1.0
0.60 2.85 3.55 3.55 3.75 1.3 3.0 4.5 3.5 3.0 1.0 3.0 4.7 3.5 2.8 1.0
0.50 3.2 3.9 3.35 3.45 1.1 3.2 4.4 3.5 2.9 1.0 3.0 4.7 3.5 2.8 1.0
0.40 3.25 3.95 3.2 3.5 1.1 3.3 4.8 3.2 2.7 1.0 3.1 4.7 3.4 2.8 1.0
0.30 3.0 3.05 3.95 3.6 1.4 3.5 4.7 3.3 2.5 1.0 3.3 4.9 3.2 2.6 1.0
0.20 2.05 2.95 4.2 3.8 2.0 3.7 4.7 3.4 2.2 1.0 3.4 4.9 3.4 2.3 1.0
0.10 2.85 3.4 3.4 3.35 2.0 3.3 4.5 3.0 2.3 1.9 3.5 4.5 3.2 2.2 1.6
0.05 3.05 4.15 3.4 2.9 1.5 2.6 4.4 3.0 2.6 2.4 3.3 4.4 3.0 2.4 1.9
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\left(\text{FGSM}(\mathcal{L},\boldsymbol{x})\right)}.$ (22)
@@
and the $\epsilon$-worst-case survival curve as the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (24)
For the exponential CPH distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq. 23. We evaluate the SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ is no perturbation to input covariates, for the two perturbations methods to study how sensitive each adversarial training method is to covariate perturbation.
Note that when the activation functions are piecewise linear, an exact solution can also be obtained by solving a Mixed Integer Linear Program (MILP). However, solving the MILP is computationally prohibitive: even for $\epsilon=0.1$ each MILP takes on average 9-10 seconds when tested on the Dialysis dataset, whereas CROWN-IBP takes milliseconds to calculate over all the data in the Dialysis dataset.
@@ VI-D Results
Furthermore, we analyze how adversarial perturbations impact the characteristics of the population survival curves.
VI-D1 Failure Rate Distribution
Let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates $\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$. Moreover, we treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process :
$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}$ (25)
Furthermore, we compute statistical quantities such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
Fig. 1 visualizes the credible interval and mean of the survival function stochastic process in Eq. 25. The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 1). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 1). The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 21, we trade-off individualized survival curves with vulnerability to adversarial perturbations and a population survival curve which is completely resilient to adversarial perturbations.
However, despite exhibiting significantly higher adversarial robustness, the credible interval usually remains relatively unchanged compared to the DRAFT method.
VI-D2 Population Survival Curve

### v4

@@ IV-B1 Fast Gradient Sign Method (FGSM)
FGSM [54] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [13] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation of $\boldsymbol{x}$ to remain within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
@@
$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),\forall k$
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)}$
where $k\in[1,\ldots,K-1]$ is the iteration step and $\alpha=\epsilon/K$ is a step size parameter.
IV-B2 Random Noise
@@ V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing models robust to dataset uncertainty, generalizable, and calibrated. We propose a training objective for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and generalization on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with adversarial regularization as a Min-Max optimization problem (Eq. 12) [51, 50], as detailed in the following sections.
@@ V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise in the logarithm of measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [21]. Instead, the proposed SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of Eq. 10:
@@ VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (Section VI-A) to quantify the impact of perturbation after the robust training is complete. To demonstrate that SAWAR produces SA models that are both high-performing and robust to dataset uncertainty, this section quantifies improvements in predictive accuracy, calibration, and population curve estimation compared to common adversarial training techniques (Section IV-B) including FGSM, Random Noise and PGD, as well as SOTA NN-based SA models, such as DRAFT and AAE-Cox.
VI-A Datasets
@@
We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively.
VI-B Experiment Setup
The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM. The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions. For AAE-Cox, we adopt the publicly available code and hyperparameters [37]. For DRAFT, we use the same hyperparameters as SAWAR as described below, excluding hyperparameters related to adversarial regularization.
The NN parameters are optimized by solving Eq. (19) using stochastic gradient descent with the ADAM optimizer [67]. We set $w=\frac{1}{\text{batch\_size}}$ and use early stopping [68] with respect to the validation set to prevent overfitting. We introduce the covariate perturbations after a set number of epochs that varies per dataset. A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after the aforementioned set number of epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude of $0.5$ is achieved.
As evaluation metrics, we consider the CI [49], the IBS [69] and the NegLL at varying $\epsilon$-perturbation magnitudes from 0 to 1. We use lifelines [70] and sksurv [48] for computing SA metrics.
@@ VI-C Adversary Setup
We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B1) and a novel worst-case perturbation. The perturbed covariates $\boldsymbol{\tilde{x}}_{FGSM}$ with respect to the input $\boldsymbol{x}$ are found via Eq. 15 and the corresponding failure rate becomes:
$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\boldsymbol{\tilde{x}}_{FGSM}}.$ (20)
@@
which we solve via CROWN-IBP (Section IV-B1) by optimizing a tight approximation of the upper bound of the maximization term. Accordingly, the $\epsilon$-worst-case survival curve has the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
$\tilde{S}(t|\boldsymbol{\tilde{x}})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (22)
DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR
1.00 3.8 4.4 3.1 2.7 3.8 3.2 5.4 5.6 3.2 2.3 2.6 1.9 5.1 5.8 3.6 2.6 2.5 1.4
0.90 3.8 4.4 3.25 2.6 3.9 3.05 5.4 5.6 3.4 2.4 2.6 1.6 5.1 5.8 3.6 2.6 2.5 1.4
0.80 3.9 4.4 3.65 2.9 3.6 2.55 5.4 5.5 3.6 2.5 2.5 1.5 5.1 5.8 3.6 2.6 2.5 1.4
0.70 4.45 4.75 3.5 3.05 3.1 2.15 5.4 5.5 3.6 2.5 2.6 1.4 5.1 5.8 3.6 2.7 2.5 1.3
0.60 4.9 5.0 3.5 3.0 2.8 1.8 5.3 5.6 3.6 2.7 2.5 1.3 5.2 5.7 3.6 2.7 2.5 1.3
0.50 5.0 4.95 3.6 3.05 2.8 1.6 5.3 5.5 3.7 2.7 2.5 1.3 5.2 5.7 3.9 2.6 2.3 1.3
0.40 5.3 4.7 3.7 3.3 2.6 1.4 5.3 5.5 3.9 2.7 2.3 1.3 5.3 5.6 3.9 2.6 2.3 1.3
0.30 5.1 5.2 3.7 3.0 2.6 1.4 5.3 5.5 3.9 2.7 2.3 1.3 5.1 5.7 3.9 2.8 2.2 1.3
0.20 5.5 5.1 3.9 2.7 2.3 1.5 5.3 5.5 3.9 2.9 2.1 1.3 4.9 5.8 4.0 2.9 2.1 1.3
0.10 5.2 5.2 4.0 3.2 1.9 1.5 5.0 5.7 4.1 3.0 1.9 1.3 4.5 5.9 4.2 3.1 1.9 1.4
0.05 5.4 5.0 3.9 3.3 2.0 1.4 4.6 5.6 4.3 3.4 1.6 1.5 4.5 5.9 4.3 3.3 1.7 1.3
0.00 3.3 4.3 4.5 4.1 2.5 2.3 2.7 4.2 4.7 4.7 2.5 2.2 3.0 5.4 4.9 4.2 1.8 1.7
TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better. Bold and underline indicates the best and second best method, respectively.
DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR
1.00 3.6 3.6 3.6 4.4 4.2 1.6 2.6 4.35 3.65 3.55 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
0.90 3.4 3.9 3.6 4.3 4.4 1.4 2.9 4.25 3.65 3.35 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
0.80 3.3 3.8 3.7 4.3 4.6 1.3 2.7 4.25 3.75 3.45 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
0.70 3.4 3.8 3.8 4.1 4.6 1.3 2.8 4.65 3.65 3.15 5.75 1.0 3.1 4.8 3.5 2.7 5.9 1.0
0.60 3.2 3.65 3.5 4.2 5.3 1.15 2.8 4.95 3.8 2.7 5.75 1.0 3.1 4.8 3.5 2.7 5.9 1.0
0.50 3.0 3.6 3.5 4.1 5.6 1.2 3.1 4.8 3.8 2.7 5.6 1.0 3.1 4.8 3.5 2.7 5.9 1.0
0.40 2.9 3.8 3.4 4.0 5.8 1.1 3.3 4.8 3.7 2.5 5.7 1.0 3.1 4.8 3.6 2.6 5.9 1.0
0.30 2.8 3.9 3.6 3.3 6.0 1.4 3.6 4.6 3.5 2.4 5.9 1.0 3.1 4.9 3.6 2.5 5.9 1.0
0.20 3.2 3.6 3.9 2.9 6.0 1.4 3.7 4.9 3.4 2.3 5.6 1.1 3.5 5.1 3.3 2.3 5.8 1.0
0.10 3.2 4.0 4.2 3.6 4.0 2.0 3.9 5.5 3.6 2.9 3.8 1.3 3.9 5.8 4.1 3.1 3.0 1.1
0.05 3.4 4.2 4.4 4.1 2.6 2.3 3.6 5.1 4.3 4.1 2.2 1.7 3.7 5.6 4.4 3.9 1.9 1.5
0.00 3.3 4.3 4.5 4.1 2.5 2.3 2.7 4.2 4.7 4.7 2.5 2.2 3.0 5.4 4.9 4.2 1.8 1.7
TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better. Bold and underline indicates the best and second best method, respectively.
For the exponential CPH distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We evaluate SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ represents no perturbation to input covariates, in order to study how sensitive different adversarial training methods are to covariate perturbations.
To simulate realistic conditions, we assume that the 10 benchmark survival analysis datasets are “clean” and contain various covariates (e.g., age, drug dosage in mg, tumor size) and survival outcomes (e.g., time until death, censored observations). By applying adversarial attack techniques, we perturb these input covariates to maximize the survival analysis error.
The perturbations are designed to be subtle yet impactful, meaning they are small in magnitude but cause significant changes in the predicted survival times. For example, since the data is standard normalized, a perturbation of $\epsilon=0.1$ would shift an input covariate by 0.1 standard deviations from its current value.
For instance, consider a scenario where one of the covariates is the body weight of a patient in kilograms. If patient $i$ has a body weight of 70 kg, an adversarial perturbation might increase this body weight to to 75 kg. While this change is minimal, it is selected because it may cause a drastic change in the predicted survival likelihood over the next 10 years for a non-robust NN-based SA model. In reality, a 5 kg change in weight should not have such a significant effect on survival, but the model may incorrectly interpret this small perturbation as a substantial change in survival hazard, highlighting its sensitivity to minor changes in input covariates.
@@ VI-D Results
We analyze how adversarial perturbations impact the characteristics of the population survival curves.
VI-D1 Comparison to SOTA
We show the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations in Tables III and II, respectively. We empirically show that SAWAR through CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes. Thus, SAWAR has better generalization performance, better calibration, and better adversarial robustness by up to 150% compared to baseline adversarial training methods and SOTA NN-based SA models. More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case and the FGSM adversarial attacks for each dataset are available in the code repository linked in the abstract.
VI-D2 Failure Rate Distribution
Each instance $x$ in the dataset has its own survival curve $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$. All instance’s survival curves are visualized as the credible interval, where $x$ is sampled from the dataset distribution $p(x)$. Thus, let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates $\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$. Moreover, we treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process :
$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}.$ (23)
We compute the statistical quantities of $S_{\boldsymbol{\theta}}(t)$ such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
Fig. 2 visualizes the credible interval and mean of the survival function stochastic process in Eq. 23. The 90% credible interval of the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 2). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 2). However, this trend has an exception with AAE-Cox. We hypothesize that since AAE-Cox is not truly an adversarial robust method (i.e., no perturbations to the input covariates during training), and by introducing more parameters via an encoder of a GAN, there are more model parameters to attack. The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 19, we introduce a trade-off between individualized survival curves with vulnerability to adversarial perturbations and a population survival curve that is completely resilient to adversarial perturbations. In doing so, while exhibiting significantly higher adversarial robustness, SAWAR’s credible interval remains relatively unchanged compared to the DRAFT method.
VI-D3 Population Survival Curve